Adv. Math. Econ. 6, 69-83 (2004)

Approximation of expectation of diffusion processes based on Lie algebra and Malliavin calculus

Shigeo Kusuoka

Graduate School of Mathematical Sciences, The University of Tokyo, 3-8-1 Komaba, Meguro-ku, Tokyo 153-8914, Japan

Received: July 9, 2003 Revised: September 17, 2003 JEL classification: C63, G12 Mathematical Subject Classification (2000): 65C05, 60G40 Abstract. The author gives a new numerical computation method of Expectation of Diffusion Processes, which is an improvement of a results in [3]. Key words: mathematical finance, option pricing, Lie algebra